9. (a) Let X be a binomial variate with n =100, p = .1.0 Find the approximate value of

P 10( ≤ X ≤12) using: (5) 

 (i) normal distribution 

 (ii)poisson distribution 

 [You may like to use the following values. 

 P(Z ≤ 67.0 ) = .0 7486, P(Z ≤ 33.0 ) = .0 6293, P(Z ≤ )0 = ]5.0

 


1
Expert's answer
2021-03-29T15:43:01-0400

i)

"\mu=np=100\cdot0.1=10"

"\sigma=\sqrt{np(1-p)}=\sqrt{100\cdot0.1(1-0.1)}=3"

"P(10 \leq X \leq12)\approx P(9.5 \leq Y \leq12.5)="

"=P(\frac{9.5-10}{3} \leq Z\leq\frac{12.5-10}{3})=P(-0.17\leq Z\leq0.83)="

"=0.7967-0.4325=0.3642"


ii)

"\lambda=np=10"

"P(X=k)=\frac{\lambda^ke^{-\lambda}}{k!}"

"P(10 \leq X \leq12)=P(X=10)+P(X=11)+P(X=12)"

"P(X=10)=\frac{10^{10}e^{-10}}{10!}=0.1251"

"P(X=11)=\frac{10^{11}e^{-10}}{11!}=0.1137"

"P(X=12)=\frac{10^{12}e^{-10}}{12!}=0.0948"

"P(10 \leq X \leq12)=0.1251+0.1137+0.0948=0.3336"